### v1

@@ 2.2 Neural ODEs for Latent Dynamics
where $n_{t}$ corresponds to the so-called training trajectory length, and $N_{T}$ represents the total number of ground-truth latent trajectories obtained at a respective $n_{t}$. In this work, the ground-truth latent trajectories in Eq. 4 are obtained from an the action of a convolutional neural network (CNN) based encoder on the corresponding full-order trajectory (i.e., ${\cal T}_{w,i}\leftarrow\text{Encode}({\cal T}_{u,i})$, where ${\cal T}_{u,i}$ contains the full-order snapshots).
Neural ODEs leverage the above trajectory data to learn a continuous-time model for the unknown latent dynamical system that governs the evolution of ${\bf w}$ chen2018neural . The starting point is to cast the functional form for the latent dynamics (the right-hand-side of Eq. 2) as a neural network, resulting in the neural ODE
@@
Figure 2: (a) Schematic of neural ODE operation. The architecture of ${\cal N}$ is a feed-forward neural network with four hidden layers and an ELU activation functions. The number of hidden units is 120 neurons. (b) Schematic of encoder operation. Encoder contains a sequence of 1D convolutional layers with batch normalization and an ELU activation function, in which the physical space input ${\bf u}(t)$ is progressively down-sampled in space. The spatial component is down-sampled by a factor of two while doubling the number of channels, upon encountering a flattening operation and linear layer to produce the latent space projection ${\bf w}(t)$. Although not shown, the decoder $\psi$ is a mirrored version of the encoder, with transpose convolution layers replacing convolution layers. Both architectures are implemented in PyTorch pytorch .
@@
where the angled brackets $\langle$.$\rangle$ . ⟩ represent an average over a batch of training set trajectories. The formulation in Eq. 7 reveals the primary advantage of the neural ODE formulation, in that the training approach is dynamics-informed: the latent dynamical system is trained to minimize accumulated error in the latent trajectory over all $n_{t}$ time steps. As such, it must be emphasized that $n_{t}$ – the training trajectory length – is a critical training parameter that captures the amount of dynamical information used to construct the latent dynamical system, and is typically chosen and fixed a-priori. The implication is that the training procedure can be executed with different $n_{t}$ values from the same time-ordered data, resulting in neural ODEs with different predictive capabilities and stability properties. To illustrate this concept, the schematic in Fig. 3 shows how a single large ground-truth latent trajectory can be split into $N_{T}$ training trajectories, each of length $n_{t}$. A primary goal of this work is to rigorously study the effect of $n_{t}$ on both predictive accuracy and latent time-scales, in addition to proposing a pathway to choose an optimal value of $n_{t}$ based on the underlying physics of the problem.
@@
It should be noted that in the training stage, in order to calculate the derivatives necessary for optimization ( $\frac{\partial\mathcal{L}_{\text{node}}}{\partial\theta}$) two approaches can be used: one involves storing the entire rollout graph of the integrated ODE trajectory and propagating gradients backward through it, while the other entails solving a system of ODEs known as the adjoint equations backward in time. In this study, the latter method is used to ensure memory efficiency when dealing with long training trajectory lengths. This work leverages the torchdiffeq library torchdiffeq for neural ODE training routines.
@@ 2.3 Dimensionality Reduction with Autoencoders
The objective of the decoder $\psi$ is to then reconstruct the original state from this latent representation with minimal loss of information, i.e. ${\bf u}(t)\approx\widetilde{\bf u}(t)=\psi(\phi({\bf u}(t)))$. Inspired by recent work in data-based reduced order modeling romit_cnn_advection ; Wan2023 ; lee2020model , this work leverages convolutional neural networks as the backbone for encoder and decoder architectures. Figure 2(right) illustrates the encoder architecture; the decoder is a mirrored version of the encoder, which transpose convolution layers replacing convolution layers. Alongside the architectural configuration, the critical parameter in any autoencoder is the latent space dimensionality $N_{w}$, which effectively controls the trade-off between dimensionality reduction and reconstruction accuracy.
@@ 2.5 Evaluation Metrics
$\mathcal{R}_{\text{RO}}=\Biggl{\langle}\sum^{n_{RO}}_{j=1}\frac{\left|\psi(\mathbf{\widetilde{{w}}}(t+\Delta t))-\mathbf{u}(t+j\Delta t)\right|}{\left|\mathbf{u}(t+j\Delta t)\right|}\Biggr{\rangle}.$ (14)
@@ 2.6 Dynamical System Timescales
Eigenvalue analysis has been used to analyse the time-scales of dynamical systems Maas1992 ; Mease2003 ; Valorani2001 . Training neural ODEs on long trajectories is a challenging problem which requires a lot of fine-tuning. The choice of the training trajectory length $n_{t}$ is often based on heuristics and does not have any theoretical justification. The complexity of the training loss function grows with the increase in $n_{t}$, and a neural ODE trained on a small $n_{t}$ may not capture the relevant dynamics of the system. Thus, choosing an optimal $n_{t}$ is an open research problem. In this work, we employ an eigenvalue based time-scale analysis framework to rigorously study the effect of $n_{t}$ on accuracy and time-scale reduction in the latent space. We employ the same framework to study the effect of certain network architecture based hyperparamaters on time-scale reduction in the latent space. For a dynamical system that is governed by an ODE given by Eq. (2), the inverse of the largest eigenvalue of the right-hand side Jacobian,
@@ 3 Numerical Experiments
[…]
@@ 3.1 KS Equation
The KS equation is solved using jax-cfd Kochkov_2021 , which allows differentiation through the full-order (physical space) dynamical system $d{\bf u}/dt=\mathbf{F(u)}$ , which facilitates Jacobian computation $\frac{\partial\mathbf{F}}{\partial\mathbf{u}}$ for extracting the corresponding true physical space time-scales facilitating comparison with the time-scales in the latent space. The KS equation is defined by the partial differential equation,
@@
where $\omega_{j}$ is randomly chosen from $\{\pi/L,2\pi/L,3\pi/L\}$, $a_{j}$ is sampled from a uniform distribution in $[-0.5,0.5]$, and the phase $\phi_{j}$ follows a uniform distribution in $[0,2\pi]$. The parameter $n_{c}=30$ governs the number of modes in the initial conditions. $\nu=0.01$ and other mentioned parameters are chosen to match previous work Wan2023 ; dresdner2023learning ; Bar_Sinai_2019 .
@@
Figure 4 compares the ground truth field for $u$, with the rollout field ($n_{RO}=500$) predicted by the autoencoder neural ODE framework trained in a decoupled manned using $n_{t}=500$, for an unseen initial condition. The predictions closely match the ground truth data. The figure also illustrates the corresponding latent space trajectories for $n_{t}=500$ and $n_{t}=4000$, both of which appear smoother compared to the state space fields. Although the presence of smooth latent space trajectories has been empirically observed Wan2023 , it has not been quantitatively studied. It can be seen that the latent space trajectories predicted by the neural ODE with $n_{t}=500$ more closely matches the ground truth trajectories as compared to the neural ODE with $n_{t}=4000$. These noticeable differences in accuracy suggest that the relationship between increasing $n_{t}$ and predictive strength is not straightforward, as discussed later. Additionally, latent space trajectories are smoother than state space fields, with $n_{t}=4000$ trajectories being smoother than $n_{t}=500$.
@@ 3.1.1 Effect of network hyperparameters
[…]
@@
[…]
@@ 3.2 Extension to Channel Detonations
In this section, to further demonstrate the generality of the timescale relationships described in the context of the KS equations in the previous sections, additional studies of the neural ODE strategy are performed on the compressible reacting Navier-Stokes equations. More specifically, surrogate models are constructed for unsteady gaseous detonations in a one-dimensional channel configuration, constituting a more realistic and highly advection-dominated demonstration case that incorporates multiple transported variables. In broad terms, an unsteady detonation can be interpreted as a propagating shockwave coupled with a chemical reaction zone shepherd . Accelerated simulations of detonation-containing flows are of significant interest to the high-speed propulsion community, where emerging concepts reliant on detonation-based combustion offer pathways for higher efficiency and robust designs shepherd ; venkat_arfm_2023 . In this context, similar to the KS demonstration studies in the previous sections, numerical solutions of the compressible NS equations with detailed chemical kinetics are used to generate ground-truth trajectory data describing the propagation of self-sustained detonation waves, from which the combined autoencoder-neural ODE models are trained. The objective is to not only showcase the capability of neural-ODE based approaches in modeling detonation dynamics, but also to illustrate consistency in neural ODE timescale trends (namely, the relationship between $n_{t}$ and the latent timescales) across fundamentally different PDEs.
@@
The detonation configuration and initiation strategy is shown in Fig. 12. More specifically, unsteady Hydrogen-Air detonations are initialized in the manner of Ref. yungster2004pulsating , where a driver gas at elevated pressure and temperature near the left wall is used to establish a self-sustained Chapman-Jouguet detonation wave that propagates through the channel. To generate ground-truth detonation data to train the data-based models, the compressible reacting NS equations are solved using an in-house flow solver developed at the University of Michigan based on the AMReX framework amrex . The solver is a block-structured adaptive mesh refinement (AMR) extension of the extensively verified UMReactingFlow umreactingflow (note that although an AMR-based solver is employed here, grid refinement is not used in this study). A globally second-order finite-volume strategy is utilized, where advection terms are treated with slope-limited Harten-Lax-van Leer-Contact approximate Riemann solver batten_1997 and diffusion terms are treated using standard central schemes. Detailed chemistry kinetics routines (species production rate and transport coefficient evaluations) are handled by Cantera cantera . In this work, hydrogen-air chemistry is modeled using the 9 species, 21 reaction detailed mechanism of Mueller et al. (1999) hon_mueller . Global time integration is handled using a Strang splitting strategy; chemical time integration is performed using an adaptive explicit method, and the advection-diffusion temporal advance uses a stability-preserving second-order Runge-Kutta method. The reader is directed to Ref. umreactingflow for additional detail on the solver numerics and discretization approach.
@@
Figure 12: Schematic of channel detonation configuration and initial condition, where a high-energy driver gas is used to initialize a self-sustained detonation wave that propagates left-to-right through the channel. Ambient and driver pressures for the trajectories considered are provided in Table 1. Driver gas compositions (not shown in schematic) come from Chapman-Jouguet conditions obtained from the Shock and Detonation toolbox sdtoolbox . Initial fluid velocity is zero throughout entire domain.
@@
For all ground-truth simulation trajectories, the detonation channel length was set to $0.3$ m, the grid resolution was fixed to 50 $\mu$m. This resolution was sufficient enough to allow for proper formation of a self-sustained detonation, allowing all complex reacting flow dynamics to be captured. The simulation time step was fixed to $5$ ns, and snapshots were written at intervals of $\Delta t=10^{-8}$ for neural ODE training and inference. Examples of detonation wave evolution, along with the corresponding smooth latent space trajectories are shown in Figure 13. Figure 14 shows the peak pressure traces for both the considered trajectories.
@@ 3.2.2 Effect of network hyperparameters
Figure 18 illustrates the limiting time-scale in the latent space, denoted as $t_{lim}$, over time for neural ODEs trained with different values of $n_{t}$. The graph indicates that increasing $n_{t}$ leads to higher values for $t_{lim}$, resulting in smoother latent space trajectories. This trend aligns with the observations for the KS equations shown in Fig. 8. Comparing the latent space’s limiting ($t_{lim}(t)$) and largest ($t_{max}(t)$) time-scales to their physical space counterparts, as in the KS equation, poses a challenge because the code generating the data prevents differentiation through the full-order dynamical system, essential for calculating Jacobians ($\frac{\partial F}{\partial u}$) required to determine the full-system time-scales.

### v3

@@ 2.2 Neural ODEs for Latent Dynamics
where $n_{t}$ corresponds to the so-called training trajectory length, and $N_{T}$ represents the total number of ground-truth latent trajectories obtained at a respective $n_{t}$. In this work, the ground-truth latent trajectories in Eq. 4 are obtained from the action of a convolutional neural network (CNN) based encoder on the corresponding full-order trajectory (i.e., ${\cal T}_{w,i}\leftarrow\text{Encode}({\cal T}_{u,i})$, where ${\cal T}_{u,i}$ contains the full-order snapshots).
Neural ODEs leverage the above trajectory data to learn a continuous-time model for the unknown latent dynamical system that governs the evolution of ${\bf w}$ [46]. The starting point is to cast the functional form for the latent dynamics (the right-hand-side of Eq. 2) as a neural network, resulting in the neural ODE
@@
Figure 2: (a) Schematic of neural ODE operation. The architecture of ${\cal N}$ is a feed-forward neural network with four hidden layers and an ELU activation functions. The number of hidden units is 120 neurons. (b) Schematic of encoder operation. Encoder contains a sequence of 1D convolutional layers with batch normalization and an ELU activation function, in which the physical space input ${\bf u}(t)$ is progressively down-sampled in space. The spatial component is down-sampled by a factor of two while doubling the number of channels, upon encountering a flattening operation and linear layer to produce the latent space projection ${\bf w}(t)$. Although not shown, the decoder $\psi$ is a mirrored version of the encoder, with transpose convolution layers replacing convolution layers. Both architectures are implemented in PyTorch [66].
@@
where the angled brackets $\langle$.$\rangle$ . ⟩ represent an average over a batch of training set trajectories. The formulation in Eq. 7 reveals the primary advantage of the neural ODE formulation, in that the training approach is dynamics-informed: the latent dynamical system is trained to minimize accumulated error in the latent trajectory over all $n_{t}$ time-steps. As such, it must be emphasized that $n_{t}$ – the training trajectory length – is a critical training parameter that captures the amount of dynamical information used to construct the latent dynamical system, and is typically chosen and fixed a-priori. The implication is that the training procedure can be executed with different $n_{t}$ values from the same time-ordered data, resulting in neural ODEs with different predictive capabilities and stability properties. To illustrate this concept, the schematic in Fig. 3 shows how a single large ground-truth latent trajectory can be split into $N_{T}$ training trajectories, each of length $n_{t}$. A primary goal of this work is to rigorously study the effect of $n_{t}$ on both predictive accuracy and latent time-scales, in addition to proposing a pathway to choose an optimal value of $n_{t}$ based on the underlying physics of the problem.
@@
It should be noted that in the training stage, in order to calculate the derivatives necessary for optimization ( $\frac{\partial\mathcal{L}_{\text{node}}}{\partial\theta}$) two approaches can be used: one involves storing the entire rollout graph of the integrated ODE trajectory and propagating gradients backward through it, while the other entails solving a system of ODEs known as the adjoint equations backward in time. In this study, the latter method is used to ensure memory efficiency when dealing with long training trajectory lengths. This work leverages the torchdiffeq library [67] for neural ODE training routines.
@@ 2.3 Dimensionality Reduction with Autoencoders
The objective of the decoder $\psi$ is to then reconstruct the original state from this latent representation with minimal loss of information, i.e. ${\bf u}(t)\approx\widetilde{\bf u}(t)=\psi(\phi({\bf u}(t)))$. Inspired by recent work in data-based reduced order modeling [27, 55, 28], this work leverages convolutional neural networks as the backbone for encoder and decoder architectures. Figure 2(right) illustrates the encoder architecture; the decoder is a mirrored version of the encoder, which transpose convolution layers replacing convolution layers. Alongside the architectural configuration, the critical parameter in any autoencoder is the latent space dimensionality $N_{w}$, which effectively controls the trade-off between dimensionality reduction and reconstruction accuracy.
@@ 2.5 Evaluation Metrics
$\mathcal{R}_{\text{RO}}=\Biggl{\langle}\sum^{n_{RO}}_{j=1}\frac{\left|\psi(\mathbf{\widetilde{{w}}}(t+j\Delta t))-\mathbf{u}(t+j\Delta t)\right|}{\left|\mathbf{u}(t+j\Delta t)\right|}\Biggr{\rangle}.$ (14)
@@ 2.6 Dynamical System Timescales
Eigenvalue analysis has been used to analyse the time-scales of dynamical systems [68, 69, 70]. Training neural ODEs on long trajectories is a challenging problem which requires a lot of fine-tuning. The choice of the training trajectory length $n_{t}$ is often based on heuristics and does not have any theoretical justification. The complexity of the training loss function grows with the increase in $n_{t}$, and a neural ODE trained on a small $n_{t}$ may not capture the relevant dynamics of the system. Thus, choosing an optimal $n_{t}$ is an open research problem. In this work, we employ an eigenvalue based time-scale analysis framework to rigorously study the effect of $n_{t}$ on accuracy and time-scale reduction in the latent space. We employ the same framework to study the effect of certain network architecture based hyperparamaters on time-scale reduction in the latent space. For a dynamical system that is governed by an ODE given by Eq. (2), the inverse of the largest eigenvalue of the right-hand side Jacobian,
@@ 3 Numerical Experiments
To assess the impact of the number of convolutional layers, latent dimensions in the autoencoder, and the training methodology, networks with varied configurations for each parameter (while keeping the other two fixed) are trained. The resulting $t_{max}(\textit{nODE},t)$ profiles are then compared to the full system time-scales provided by Eq.(15). Furthermore, the influence of the training trajectory length $n_{t}$ on both time-scale reduction and predictive accuracy is considered for both numerical experiments.
@@ 3.1 KS Equation
The KS equation is solved using jax-cfd [71], which allows differentiation through the full-order (physical space) dynamical system $d{\bf u}/dt=\mathbf{F(u)}$ , which facilitates Jacobian computation $\frac{\partial\mathbf{F}}{\partial\mathbf{u}}$ for extracting the corresponding true physical space time-scales facilitating comparison with the time-scales in the latent space. The KS equation is defined by the partial differential equation,
@@
where $\omega_{j}$ is randomly chosen from $\{\pi/L,2\pi/L,3\pi/L\}$, $a_{j}$ is sampled from a uniform distribution in $[-0.5,0.5]$, and the phase $\phi_{j}$ follows a uniform distribution in $[0,2\pi]$. The parameter $n_{c}=30$ governs the number of modes in the initial conditions. $\nu=0.01$ and other mentioned parameters are chosen to match previous work [55, 72, 73].
@@
Figure 4 compares the ground truth field for $u$, with the rollout field ($n_{RO}=500$) predicted by the autoencoder neural ODE framework trained in a decoupled manner using $n_{t}=500$, for an unseen initial condition. The predictions closely match the ground truth data. The figure also illustrates the corresponding latent space trajectories for $n_{t}=500$ and $n_{t}=4000$, both of which appear smoother compared to the state space fields. Although the presence of smooth latent space trajectories has been empirically observed [55], it has not been quantitatively studied. It can be seen that the latent space trajectories predicted by the neural ODE with $n_{t}=500$ more closely matches the ground truth trajectories as compared to the neural ODE with $n_{t}=4000$. These noticeable differences in accuracy suggest that the relationship between increasing $n_{t}$ and predictive strength is not straightforward, as discussed later. Additionally, latent space trajectories are smoother than state space fields, with $n_{t}=4000$ trajectories being smoother than $n_{t}=500$.
@@ 3.1.1 Effect of network hyperparameters
However, Fig. 5 illustrates the limiting time-scale in the latent space, denoted as $t_{lim}(\textit{nODE},t)$, as a function of time for autoencoders with varying numbers of convolutional layers. Contrary to expectations, the number of convolutional layers does not exert a significant influence on the time-scales of the latent variables. In Fig. 5, the line corresponding to the ’full-system’ is generated by plotting the inverse of the largest eigenvalue of the full KSE-equations right-hand side Jacobian as a function of time. It can be seen that all the autoencoders in general significantly reduce the limiting time-scales of the system, by roughly six orders of magnitude.
@@
Latent space dimensionality: Figure 5 also depicts the limiting time-scale in the latent space, denoted as $t_{lim}(\textit{nODE},t)$, as a function of time for autoencoders with different numbers of latent dimensions. The observation from the figure is that the size of the latent space does not have a substantial impact on the limiting time-scales in the latent space.
@@ 3.2 Extension to Channel Detonations
In this section, to further demonstrate the generality of the timescale relationships described in the context of the KS equations in the previous sections, additional studies of the neural ODE strategy are performed on the compressible reacting Navier-Stokes equations. More specifically, surrogate models are constructed for unsteady gaseous detonations in a one-dimensional channel configuration, constituting a more realistic and highly advection-dominated demonstration case that incorporates multiple transported variables. In broad terms, an unsteady detonation can be interpreted as a propagating shockwave coupled with a chemical reaction zone [74]. Accelerated simulations of detonation-containing flows are of significant interest to the high-speed propulsion community, where emerging concepts reliant on detonation-based combustion offer pathways for higher efficiency and robust designs [74, 75]. In this context, similar to the KS demonstration studies in the previous sections, numerical solutions of the compressible NS equations with detailed chemical kinetics are used to generate ground-truth trajectory data describing the propagation of self-sustained detonation waves, from which the combined autoencoder-neural ODE models are trained. The objective is to not only showcase the capability of neural-ODE based approaches in modeling detonation dynamics, but also to illustrate consistency in neural ODE timescale trends (namely, the relationship between $n_{t}$ and the latent timescales) across fundamentally different PDEs.
@@
The detonation configuration and initiation strategy is shown in Fig. 12. More specifically, unsteady Hydrogen-Air detonations are initialized in the manner of Ref. [76], where a driver gas at elevated pressure and temperature near the left wall is used to establish a self-sustained Chapman-Jouguet detonation wave that propagates through the channel. To generate ground-truth detonation data to train the data-based models, the compressible reacting NS equations are solved using a flow solver developed at the University of Michigan based on the AMReX framework [77, 78]. The solver is a block-structured adaptive mesh refinement (AMR) extension of the extensively verified UMReactingFlow [79] (note that although an AMR-based solver is employed here, grid refinement is not used in this study). A globally second-order finite-volume strategy is utilized, where advection terms are treated with slope-limited Harten-Lax-van Leer-Contact approximate Riemann solver [80] and diffusion terms are treated using standard central schemes. Detailed chemistry kinetics routines (species production rate and transport coefficient evaluations) are handled by Cantera [81]. In this work, hydrogen-air chemistry is modeled using the 9 species, 21 reaction detailed mechanism of Mueller et al. (1999) [82]. Global time integration is handled using a Strang splitting strategy; chemical time integration is performed using an adaptive explicit method, and the advection-diffusion temporal advance uses a stability-preserving second-order Runge-Kutta method. The reader is directed to Ref. [79] for additional detail on the solver numerics and discretization approach.
@@
Figure 12: Schematic of channel detonation configuration and initial condition, where a high-energy driver gas is used to initialize a self-sustained detonation wave that propagates left-to-right through the channel. Ambient and driver pressures for the trajectories considered are provided in Table 1. Driver gas compositions (not shown in schematic) come from Chapman-Jouguet conditions obtained from the Shock and Detonation toolbox [83]. Initial fluid velocity is zero throughout entire domain.
@@
For all ground-truth simulation trajectories, the detonation channel length was set to $0.3$ m, the grid resolution was fixed to 50 $\mu$m. This resolution was sufficient enough to allow for proper formation of a self-sustained detonation, allowing all complex reacting flow dynamics to be captured. The simulation time-step was fixed to $5$ ns, and snapshots were written at intervals of $\Delta t=10^{-8}$ for neural ODE training and inference. Examples of detonation wave evolution, along with the corresponding smooth latent space trajectories are shown in Figure 13. Figure 14 shows the peak pressure traces for both the considered trajectories.
@@ 3.2.2 Effect of network hyperparameters
Figure 18 illustrates the limiting time-scale in the latent space, denoted as $t_{lim}$, over time for neural ODEs trained with different values of $n_{t}$. The graph indicates that increasing $n_{t}$ leads to higher values for $t_{lim}$, resulting in smoother latent space trajectories. This trend aligns with the observations for the KS equations shown in Figure. 8. Comparing the latent space’s limiting ($t_{lim}(t)$) and largest ($t_{max}(t)$) time-scales to their physical space counterparts, as in the KS equation, poses a challenge because the code generating the data prevents differentiation through the full-order dynamical system, essential for calculating Jacobians ($\frac{\partial F}{\partial u}$) required to determine the full-system time-scales.
@@
3.3 Extension to 2D Atmospheric flow
In this section, additional demonstrations on an atmospheric flow dataset are performed to augment the results from the KS equation and detonation test cases, and to further verify trends with respect to the training trajectory length on the neural ODE strategy. Similar to Sec. 3.2, the details of the dataset are first provided, followed by an assessment of model performance and $n_{t}$ effects. Additionally, this section also adds to the preexisting analysis above, in that the effect of simulation time-step variations on neural ODE performance is also assessed.
The 2D atmospheric fields are produced using a simplified atmospheric general circulation model (AGCM) called SPEEDY. SPEEDY is a spectral transform AGCM that was developed to produce rapid climate simulations, using simplified, but modern physical parameterization schemes [84, 85]. The grid of SPEEDY consists of a horizontal spatial resolution of ${}^{\circ}\times$3.75 with variables defined at eight vertical levels. The 3D varying state variables of the model are the two components of the horizontal wind vector, temperature, and specific humidity (moisture), while the single two-dimensionally varying state variable is the natural logarithm of surface pressure. For this dataset, we subsample the 3D variables to obtain 4 key variables of the model: near surface temperature, specific humidity, and upper level winds.
This atmospheric dataset constitutes a highly complex and nonlinear advection-dominant benchmark that is fundamentally different from both KSE and detonation cases described above. It therefore serves as an additional dynamical system in this work to more comprehensively assess neural ODE timescale trends. For additional details on the physical phenomena contained in the dataset, the reader is directed to Refs. [84, 85].
3.3.1 Extrapolation in time
The predictive performance of the trained autoencoder neural ODE framework, developed using the decoupled approach and configured with $n_{t}=20$, $n_{\hat{u}}=200$, and four convolutional layers in the autoencoder (see Figure 2), is assessed by extrapolating beyond the training dataset. Figure 19 presents the predicted and ground truth fields for temperature, specific humidity (SH), and the $u$ and $v$ velocity components at various time instances, both within and beyond the training period. The results suggest that the neural ODE autoencoder framework captures the system dynamics within the training domain and shows an ability to predict system behavior at unseen time instances. It should be noted that while many of the large scale features are preserved at the shown extrapolation time, there are indeed unphysical fine-scale artifacts in the predicted solution, pointing to intrinsic instabilities for this problem. Stability challenges for this dataset are generally expected using CNN-based encodings, due to the high level of nonlinearity and the fact that no additional stability enhancements are employed here. Despite this, it is emphasized that the purpose of this additional demonstration is to augment the KS equation and detonation analysis through inspection/understanding of the role of $n_{t}$, which is summarized below.
Figure 19: Comparison of the ground truth and predicted fields for atmospheric dataset the within the training set (left) and outside the training set (right). Results shown for $n_{t}=20$, $N_{w}=200$, and four convolutional layers, with extrapolation time of 2500 time-steps)
3.3.2 Effect of network hyperparameters
Figure 20 illustrates the Mean Squared Error (MSE) of predicted trajectories beyond the training set, evaluated over 1000 samples, as a function of $n_{t}$. The plot highlights that the error reaches its minimum at an optimal $n_{t}$, consistent with trends observed in Figure 10 for the KS equations and Figure 17 for the 1D channel detonation case.
Figure 20: $\mathcal{L}_{\text{RO}}$ (Eq. 12) for a rollout trajectory length $n_{\text{RO}}=1000$, as a function of $n_{t}$ ($N_{w}=100$ and four convolutional layers).
Figure 21 shows the evolution of the limiting time-scale in the latent space, $t_{lim}$, for neural ODEs trained with varying $n_{t}$ values. The results show that larger $n_{t}$ values correspond to increased $t_{lim}$, which promotes smoother trajectories in the latent space. This behavior is consistent with the trends observed for the KS equations in Figure 8 and the 1D channel detonation case in Figure 18.
Figure 21: Limiting time-scale in the latent space ($t_{lim}$) (Eq. 17) as a function of time for neural ODEs with different training trajectory lengths $n_{t}$ ($N_{w}=100$ and four convolutional layers)
3.3.3 Effect of time-step
The effect of the time-step size $\Delta t$, used in the latent space during inference, on the rollout error $\mathcal{L}_{\text{RO}}$ is analyzed in this section. For this analysis, the best-performing model with $n_{t}=20$, trained using $\Delta t=10$, was selected.
Figure 22 illustrates the variation of the rollout error $\mathcal{L}_{\text{RO}}$ as a function of the time-step $\Delta t$ used in the latent space for a rollout trajectory length of $n_{\text{RO}}=1000$. The results indicate that the error generally decreases as the time-step increases and eventually reaches a plateau. This behavior suggests that neural ODEs trained on discrete data points may primarily approximate discrete dynamics rather than capturing the underlying continuous dynamics. As a result, when the inference time-step deviates from the training time-step, the model’s error can increase significantly. Notably, evaluating the model at different time-step sizes, whether larger or smaller than the training step, can lead to substantial errors, indicating potential overfitting to the training discretization and limited generalization across different time-step scales. Similar trends have also been observed in previous work [86, 87, 88], highlighting the challenges of generalizing across different time-step scales.
Figure 22: $\mathcal{L}_{\text{RO}}$ (Eq. 12) for a rollout trajectory length $n_{\text{RO}}=2500$, as a function of $\Delta_{t}$ ($N_{w}=100$ and four convolutional layers).